F(x) < 0 if 0<x< 2or x > 4 f(x)>0if1<x<3, F(x)<0if x<lor x>3 4. Find the limit. lim(1-2x) 10 5. What is the minimum vertic
Answer #1


Ary: clim (1-221) de ano afobly Earpomme Rule Ion (1-23) = lim í to In 11-20) a = lime 2130 apply L Hospital Rule 7*-2 -220 = det Ě o suorgno Centurile t = 1+($X$) - Chao (k) նաև lununeen ostales en 1- xtape It le cuvee : 0 (746 – )-City H) = 4,-42 D=